what is the volume of a 0.16 M sulfuric acid (H2SO4) solution that contains 0.35 mol sulfuric acid?
1
Expert's answer
2016-11-15T15:23:10-0500
C = 0.16 M
n = 0.35 mol.
V= c/n. V=0.35/0.16 = 2.22(liters)
Answer: 2.22 liters.